# READ-PVLA: Recurrent Adapter with Partial Video-Language Alignment for Parameter-Efficient Transfer Learning in Low-Resource Video-Language Modeling

### v1

@@ Abstract
Fully fine-tuning pretrained large-scale transformer models has become a popular paradigm for video-language modeling tasks, such as temporal language grounding and video-language summarization. With a growing number of tasks and limited training data, such full fine-tuning approach leads to costly model storage and unstable training. To overcome these shortcomings, we introduce lightweight adapters to the pre-trained model and only update them at fine-tuning time. However, existing adapters fail to capture intrinsic temporal relations among video frames or textual words. Moreover, they neglect the preservation of critical task-related information that flows from the raw video-language input into the adapter’s low-dimensional space. To address these issues, we first propose a novel REcurrent ADapter (READ) that employs recurrent computation to enable temporal modeling capability. Second, we propose Partial Video-Language Alignment (PVLA) objective via the use of partial optimal transport to maintain task-related information flowing into our READ modules. We validate our READ-PVLA framework through extensive experiments where READ-PVLA significantly outperforms all existing fine-tuning strategies on multiple low-resource temporal language grounding and video-language summarization benchmarks.
1 Introduction
@@
To address these shortcomings, adapters are proposed as a parameter-efficient solution for finetuning video-language pretrained transformers (Jiang et al. 2022; Yang et al. 2022; Zhang et al. 2023; Yang et al. 2023; Sung, Cho, and Bansal 2022; Chen et al. 2022). The strategy is to add additional adaptation module to each layer of the pre-trained network and only the adaptation modules are trained during fine-tuning to improve the parameter-performance trade-off. These modules rely on non-linear projections to downproject video-language inputs into low-dimensional space then up-project them back to the original high-dimensional space. However, such projections consider video frames and textual words as separate tokens, thus ignoring the intrinsic temporal dependency among video frames or textual words. Without such dependency information, it is difficult to reason about temporal context in the video to properly ground the language (e.g. in Figure 1, determine the expression of the girl after, not before, the proposal), or coherently link the entities in the summary (e.g. in Figure 1, recap the chronological order of bolt removing and puller threading). Moreover, because at fine-tuning time only adaptation modules are trained using limited video-language data, little attention is paid to the information flow that starts from the raw video-language inputs till the low-dimensional space of the adaptation modules. This may result in losing essential task-related information and carrying noise into these modules (Tsai et al. 2020; Han, Chen, and Poria 2021).
@@
Based on our novel proposals, we construct the READ-PVLA framework that can be employed to finetune various pre-trained Transformer architectures such as multimodal transformer (UMT (Liu et al. 2022), Moment-DETR (Lei, Berg, and Bansal 2021)), and generative vision-guided transformer models (VG-BART (Lewis et al. 2020) and VG-T5 (Raffel et al. 2020)). Through freezing these pre-trained models and fine-tuning only our READ modules with PVLA objective, we outperform standard fine-tuning and other parameter-efficient methods with substantially fewer tunable parameters (Figure 2) for low-resource video-language tasks, including temporal language grounding and video-language summarization. To sum up, our contributions can be summarized as:
@@
  • We validate our READ-PVLA framework by extensive experiments using multiple low-resource temporal language grounding and video-language summarization datasets, where READ-PVLA outperforms all existing fully or parameter-efficient fine-tuning strategies with only up to 1.20% parameters tunable.
2 Related Work
2.1 Parameter-Efficient Transfer Learning
Motivated by the prevalent application of pre-trained Transformer models, recent efforts have sought to propose techniques to reduce the cost of fine-tuning these large-scale models. The techniques can be categorized into three directions. The first one, dubbed as adapter, introduces lightweight modules between Transformer layers that work in low-dimensional space (Houlsby et al. 2019; Pan et al. 2022; Chen et al. 2022; Xu et al. 2023; Zhang et al. 2023; Karimi Mahabadi, Henderson, and Ruder 2021). During fine-tuning, only parameters of the adapters are updated and all of the original Transformer are kept frozen. The second approach, called prompt tuning, appends a sequence of prefix continuous tokens to every input and solely tunes these tokens for adapting to the downstream task (Li and Liang 2021; Jia et al. 2022; Huang et al. 2023). The third approach approximates the weight update with low-rank matrices (Hu et al. 2021). Only values of the matrices are learned during training to satisfy the parameter-efficiency requirement.
2.2 Video-Language Modeling
Recent video-language modeling tasks, e.g. temporal language grounding (TLG) (Liu et al. 2022; Lei, Berg, and Bansal 2021; Zeng, Pan, and Han 2023; Nguyen et al. 2023) or video-language summarization (VLS) (Yu et al. 2021; Liu et al. 2023) have been dominated by deep Transformer models. Regarding the TLG task, (Lei, Berg, and Bansal 2021) collect a query-based benchmark and construct a Transformer model pre-trained upon automatic speech recognition to tackle not only their benchmark but also other datasets. (Liu et al. 2022) follow with a multimodal Transformer that first considers video and language as separate input streams and unifies them, while preserving the pre-training scheme of (Lei, Berg, and Bansal 2021). For the VLS task, (Yu et al. 2021) take into account generative pre-trained Transformer to fuse the video and language content then generate the output summary.
@@ 3 Methodology
3.1 Preliminary - Transformer architecture for video-language modeling
@@
$\mathbf{Q}^{(m)}=\text{Linear}\left(\mathbf{H}_{L}^{(m)}\right),$ (1)
$\mathbf{K}^{(m)}=\text{Linear}\left(\mathbf{H}_{V}^{(m)}\right),$ (2)
$\mathbf{V}^{(m)}=\text{Linear}\left(\mathbf{H}_{V}^{(m)}\right),$ (3)
@@
3.2 Recurrent Adapter (READ)
@@
In addition to RNN, we also experiment with other recurrent architectures in Table 4.3 and observe that the performance is insensitive to the architectural choice. Therefore, for simplicity, we decide to implement the RNN architecture in our READ layer.
@@
3.3 Partial Video-Language Alignment (PVLA)
@@ 4 Experiments
         $\mathbf{p}_{a}=\min\left(\frac{\mathbf{a}}{\mathbf{T}\mathbf{1}_{N_{L}}},\mathbf{1}_{N_{V}}\right)$
         $\mathbf{T}_{a}=\text{diag}\left(\mathbf{p}_{a}\right)\cdot\mathbf{T}$
         $\mathbf{p}_{b}=\min\left(\frac{\mathbf{b}}{\mathbf{T}_{a}^{\top}\mathbf{1}_{N_{V}}},\mathbf{1}_{N_{L}}\right)$
         $\mathbf{T}_{b}=\text{diag}\left(\mathbf{p}_{b}\right)\cdot\mathbf{T}_{a}$
@@
We conduct extensive experiments to evaluate the effectiveness of our READ-PVLA framework. We first describe the experimental settings, covering the downstream tasks, evaluation metrics, pre-trained backbones, baseline approaches, and implementation details. We then present the numerical results of our method with baseline models, then provide ablation study and thorough analysis to explore various configurations. Eventually, we perform qualitative assessments to further elucidate the behavior of our framework.
4.1 Experimental Settings
@@
4.2 Main results
Method #params (M) Dog Gym Par. Ska. Ski. Sur. Avg.
Full 283.97 (100%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Bias 0.51 (0.18%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Partial 38.75 (13.65%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Proj 5e-4 (1.76e-4%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
LoRA 13.12 (4.62%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Prompt 0.02 (0.01%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Adapter 13.11 (4.62%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
READ-PVLA 0.16 (0.06%) 67.65 78.05 83.25 72.40 72.98 82.36 76.12
Table 1: TLG results on the YouTube Highlights dataset. We report the mean average precision (mAP) and the number of trainable parameters (#params). ${}^{\ddagger}$means the gain of READ-PVLA is statistically significant at the 0.05 level.
Method #params (M) VT VU GA MS PK PR FM BK BT DS Avg.
Full 285.28 (100%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Bias 0.25 (0.09%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Partial 38.75 (13.58%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Proj 5e-4 (1.75e-4%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
LoRA 13.28 (4.66%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Prompt 0.02 (0.007%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Adapter 13.29 (4.66%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
READ-PVLA 0.14 (0.05%) 88.30 85.15 89.76 75.80 86.69 86.62 74.99 82.38 84.65 79.60 83.39
Table 2: TLG results on the TVSum dataset. We report the mean average precision (mAP) and the number of trainable parameters (#params). ${}^{\ddagger}$means the gain of READ-PVLA is statistically significant at the 0.05 level.
For main comparison of our READ-PVLA with baseline methods, we denote the results of YouTube Highlights in Table 1, TVSum in Table 2, QVHighlights in Table 3, and How2 in Tables 4 and 5.
Temporal language grounding (TLG). For the YouTube Highlights dataset, our READ-PVLA framework substantially outperforms the Full fine-tuning approach (e.g. 1.36% on average, 1.75% in the Dog domain, and 1.21% in the Surfing domain), while updating far less parameters (0.16M vs. 283.97M). We significantly surpass all other efficient fine-tuning methods as well, e.g. with an improvement of 10.96% over the Adapter in the Gym category, or 5.78% over LoRA in the Skating category.
@@ 4.2 Main results
These results demonstrate that our READ-PVLA framework can efficiently model video-language inputs to polish the low-resource temporal language grounding performance of various pre-trained Transformer models.
Method #params (M) mAP
Full 15.88 (100%) ${}^{\ddagger}$
Bias 0.06 (0.38%) ${}^{\ddagger}$
Partial 1.05 (6.61%) ${}^{\ddagger}$
Proj 7.31 (46.03%) ${}^{\ddagger}$
LoRA 0.19 (1.20%) ${}^{\ddagger}$
Prompt 0.04 (0.25%) ${}^{\ddagger}$
Adapter 0.20 (1.26%) ${}^{\ddagger}$
READ-PVLA 0.19 (1.20%) 36.74
Table 3: TLG results on the QVHighlights dataset. We report the mean average precision (mAP) and the number of trainable parameters (#params). ${}^{\ddagger}$means the gain of READ-PVLA is statistically significant at the 0.05 level.
Method #params (M) R1 R2 RL
Full 249.67 (100%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Bias 0.20 (0.08%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Partial 16.54 (6.62%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Proj 38.60 (15.46%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
LoRA 1.19 (0.48%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Prompt 0.05 (0.02%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Adapter 1.20 (0.48%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
READ-PVLA 1.17 (0.47%) 44.01 21.91 37.91
Table 4: VLS results on the How2 dataset with the VG-BART model. We report the ROUGE-1, ROUGE-2, and ROUGE-L scores, with the number of trainable parameters (#params). ${}^{\ddagger}$means the gain of READ-PVLA is statistically significant at the 0.05 level.
Method #params (M) R1 R2 RL
Full 333.16 (100%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Bias 0.07 (0.02%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Partial 16.52 (4.96%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Proj 24.67 (7.40%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
LoRA 1.93 (0.58%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Prompt 0.18 (0.05%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Adapter 1.95 (0.59%) ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
READ-PVLA 1.91 (0.57%) 40.12 18.71 34.42
Table 5: VLS results on the How2 dataset with the VG-T5 model. We report the ROUGE-1, ROUGE-2, and ROUGE-L scores, with the number of trainable parameters (#params). ${}^{\ddagger}$means the gain of READ-PVLA is statistically significant at the 0.05 level.
@@
4.3 Ablation studies
We ablate our READ-PVLA framework to discover what factors result in the demonstrated efficiency and observe several intriguing properties. Our ablation studies are all conducted on the YouTube Highlights and How2 test set.
Method mAP - YouTube Highlights R2 - How2
No VLA 73.80 18.22
VLA 74.41 20.01
PVLA 76.12 21.91
Table 6: Partial video-language alignment (PVLA) ablation experiments on YouTube Highlights and How2. We color the settings we implement for our READ-PVLA method.
Table 7: Bottleneck dimension ablation experiments on YouTube Highlights and How2. We color the settings we implement for our READ-PVLA method.
Bottleneck dim #params - UMT (M) mAP -
YouTube Highlights
#params - VG-BART (M) R2 - How2
1
0.04 63.91 0.29 18.57
2
0.08 74.69 0.59 21.39
4
0.16 76.12 1.17 21.91
16
0.6 75.44 4.69 21.85
32
1.10 76.01 9.37 21.63
Table 8: Recurrent architecture ablation experiment on YouTube Highlights and How2. We color the settings we implement for our READ-PVLA method.
Recurrent arch. #params - UMT (M) mAP -
YouTube Highlights
#params - VG-BART (M) R2 - How2
GRU
0.16 76.07 1.17 21.85
LSTM
0.16 76.08 1.17 21.87
RNN
0.16 76.12 1.17 21.91
Position #params - UMT (M) mAP - YouTube Highlights #params - VG-BART (M) R2 - How2
1 $\rightarrow\frac{M}{2}$
0.06
75.08
0.48
21.53
$\frac{M}{2}+1\rightarrow M$
0.09
75.31
0.69
21.59
$1\rightarrow M$
0.16
76.12
1.17
21.91
Table 9: Model-level position ablation experiments on YouTube Highlights and How2. $M$ denotes the number of Transformer blocks denoted in Section 3.1. $1\rightarrow\frac{M}{2}$ means that we include Adapter from the first to the $\frac{M}{2}$ block, and similarly for $\frac{M}{2}+1\rightarrow M$. We color the settings we implement for our READ-PVLA method.
Position mAP - YouTube Highlights R2 - How2
Before Self-Attention 75.77 21.04
After Self-Attention 75.79 21.12
Before FeedForward 75.86 21.44
After FeedForward 76.12 21.91
Table 10: Block-level position ablation experiments on YouTube Highlights and How2. We color the settings we implement for our READ-PVLA method.
Distance method mAP - YouTube Highlights R2 - How2
AvgPool - Cosine 71.65 20.32
MaxPool - Cosine 74.37 21.36
AvgPool - L2 72.11 20.73
MaxPool - L2 74.39 21.02
Partial OT 76.12 21.91
Table 11: Distance method ablation experiments on YouTube Highlights and How2. We color the settings we implement for our READ-PVLA method.
Video Language query POT distance AP
[Uncaptioned image] [Uncaptioned image]
Girl’s expression after getting proposal
27.17 90.28
[Uncaptioned image] [Uncaptioned image]
A boy showing his arm after being stung at the beach
56.38 30.00
Table 12: Case study on the temporal language grounding benchmark. We extract the POT distance between video and language of two inputs with different language queries and measure the respective AP performance change. The video flows from top to bottom and left to right.

### v2

# READ: Recurrent Adapter with Partial Video-Language Alignment for Parameter-Efficient Transfer Learning in Low-Resource Video-Language Modeling

@@ Abstract
Fully fine-tuning pretrained large-scale transformer models has become a popular paradigm for video-language modeling tasks, such as temporal language grounding and video-language summarization. With a growing number of tasks and limited training data, such full fine-tuning approach leads to costly model storage and unstable training. To overcome these shortcomings, we introduce lightweight adapters to the pre-trained model and only update them at fine-tuning time. However, existing adapters fail to capture intrinsic temporal relations among video frames or textual words. Moreover, they neglect the preservation of critical task-related information that flows from the raw video-language input into the adapter’s low-dimensional space. To address these issues, we first propose a novel REcurrent ADapter (READ) that employs recurrent computation to enable temporal modeling capability. Second, we propose Partial Video-Language Alignment (PVLA) objective via the use of partial optimal transport to maintain task-related information flowing into our READ modules. We validate our READ framework through extensive experiments where READ significantly outperforms all existing fine-tuning strategies on multiple low-resource temporal language grounding and video-language summarization benchmarks. The code, model, and data have been made available at nguyentthong.github.io/READ.
Introduction
@@
To address these shortcomings, adapters are proposed as a parameter-efficient solution for finetuning video-language pretrained transformers (Jiang et al. 2022; Zhang et al. 2023; Yang et al. 2023; Sung, Cho, and Bansal 2022; Chen et al. 2022). The strategy is to add additional adaptation module to each layer of the pre-trained network and only the adaptation modules are trained during fine-tuning to improve the parameter-performance trade-off. These modules rely on non-linear projections to downproject video-language inputs into low-dimensional space then up-project them back to the original high-dimensional space. However, such projections consider video frames and textual words as separate tokens, thus ignoring the intrinsic temporal dependency among video frames or textual words. Without such dependency information, it is difficult to reason about temporal context in the video to properly ground the language (e.g. in Figure 1, determine the expression of the girl after, not before, the proposal), or coherently link the entities in the summary (e.g. in Figure 1, recap the chronological order of bolt removing and puller threading). Moreover, because at fine-tuning time only adaptation modules are trained using limited video-language data, little attention is paid to the information flow that starts from the raw video-language inputs till the low-dimensional space of the adaptation modules. This may result in losing essential task-related information and carrying noise into these modules (Tsai et al. 2020; Han, Chen, and Poria 2021).
@@
Based on our novel proposals, we construct the READ framework that can be employed to finetune various pre-trained Transformer architectures such as multimodal transformer (UMT (Liu et al. 2022), Moment-DETR (Lei, Berg, and Bansal 2021)), and generative vision-guided transformer models (VG-BART (Lewis et al. 2020) and VG-T5 (Raffel et al. 2020)). Through freezing these pre-trained models and fine-tuning only our READ modules with PVLA objective, we outperform standard fine-tuning and other parameter-efficient methods with substantially fewer tunable parameters (Figure 2) for low-resource video-language tasks, including temporal language grounding and video-language summarization. To sum up, our contributions can be summarized as:
@@
  • We validate our READ framework by extensive experiments using multiple low-resource temporal language grounding and video-language summarization datasets, where READ outperforms all existing fully or parameter-efficient fine-tuning strategies with only up to 1.20% parameters tunable.
Related Work
Parameter-Efficient Transfer Learning
Recent efforts have sought to propose techniques to reduce the cost of fine-tuning these large-scale models. The techniques can be categorized into three directions. The first one, dubbed as adapter, introduces lightweight modules between Transformer layers that work in low-dimensional space (Houlsby et al. 2019; Pan et al. 2022; Chen et al. 2022; Xu et al. 2023). During fine-tuning, only parameters of the adapters are updated and all of the original Transformer are kept frozen. The second approach, called prompt tuning, appends a sequence of prefix continuous tokens to every input and solely tunes these tokens for adapting to the downstream task (Jia et al. 2022; Huang et al. 2023). The third approach approximates the weight update with low-rank matrices (Hu et al. 2021). Only values of the matrices are learned during training to satisfy the parameter-efficiency requirement.
Video-Language Modeling
Recent video-language modeling tasks, e.g. temporal language grounding (TLG) (Liu et al. 2022; Lei, Berg, and Bansal 2021; Nguyen et al. 2023) or video-language summarization (VLS) (Yu et al. 2021; Liu et al. 2023) have been dominated by deep Transformer models. Regarding the TLG task, (Lei, Berg, and Bansal 2021) collect a query-based benchmark and construct a Transformer model pre-trained upon automatic speech recognition to tackle not only their benchmark but also other datasets. (Liu et al. 2022) follow with a multimodal Transformer that first considers video and language as separate input streams and unifies them, while preserving the pre-training scheme of (Lei, Berg, and Bansal 2021). For the VLS task, (Yu et al. 2021) take into account generative pre-trained Transformer to fuse the video and language content then generate the output summary.
@@
Preliminary – Transformer Architecture for Video-Language Modeling
@@
$\mathbf{Q}^{(m)}=\text{Linear}\left(\mathbf{H}_{L}^{(m)}\right),\;\mathbf{K}^{(m)}=\text{Linear}\left(\mathbf{H}_{V}^{(m)}\right),$
$\mathbf{V}^{(m)}=\text{Linear}\left(\mathbf{H}_{V}^{(m)}\right),$
@@
Recurrent Adapter (READ)
@@
In addition to RNN, we also experiment with other recurrent architectures in Table 8 and observe that the performance is insensitive to the architectural choice. Therefore, for simplicity, we decide to implement the RNN architecture in our READ layer.
@@
Partial Video-Language Alignment (PVLA)
@@
         $\mathbf{p}_{a}=\min\left(\frac{\mathbf{a}}{\mathbf{T}\mathbf{1}_{N_{L}}},\mathbf{1}_{N_{V}}\right)$, $\mathbf{T}_{a}=\text{diag}\left(\mathbf{p}_{a}\right)\cdot\mathbf{T}$
         $\mathbf{p}_{b}=\min\left(\frac{\mathbf{b}}{\mathbf{T}_{a}^{\top}\mathbf{1}_{N_{V}}},\mathbf{1}_{N_{L}}\right)$, $\mathbf{T}_{b}=\text{diag}\left(\mathbf{p}_{b}\right)\cdot\mathbf{T}_{a}$
@@
Method #params (M) Dog Gym Par. Ska. Ski. Sur. Avg.
Full 283.97 (100%) 65.90 75.20 82.20 71.80 72.30 81.15 74.76
Bias 0.51 (0.18%) 46.23 61.19 56.73 31.36 61.14 49.77 51.07
Partial 38.75 (13.65%) 48.28 63.26 59.71 32.66 64.58 56.22 54.12
Proj 5e-4 (1.76e-4%) 57.05 65.70 63.03 71.83 65.45 79.71 67.13
LoRA 13.12 (4.62%) 60.97 67.68 72.53 66.62 71.24 79.15 69.70
Prompt 0.02 (0.01%) 48.28 63.26 59.71 35.67 35.67 64.61 46.87
Adapter 13.11 (4.62%) 62.89 67.09 74.56 62.56 68.10 78.73 68.98
READ 0.16 (0.06%) 67.65 78.05 83.25 72.40 72.98 82.36 76.12
Table 1: TLG results on the YouTube Highlights dataset. We report the mean average precision (mAP) and the number of trainable parameters (#params). means the gain of READ is statistically significant at the 0.05 level.
Method #params (M) VT VU GA MS PK PR FM BK BT DS Avg.
Full 285.28 (100%) 84.17 81.50 88.20 71.54 81.40 84.31 72.30 76.53 78.86 77.70 79.65
Bias 0.25 (0.09%) 38.08 69.62 60.87 31.25 68.84 51.71 50.72 65.38 54.42 59.05 54.99
Partial 38.75 (13.58%) 57.27 62.57 58.08 52.35 61.58 63.94 50.82 62.36 58.05 47.79 57.48
Proj 5e-4 (1.75e-4%) 57.65 65.80 64.40 55.57 64.67 67.07 59.08 74.70 63.29 49.48 62.17
LoRA 13.28 (4.66%) 77.87 77.01 77.82 66.38 80.21 82.23 66.89 72.31 69.58 72.09 74.24
Prompt 0.02 (0.007%) 61.67 71.98 64.07 35.54 72.74 48.70 52.97 67.59 57.28 38.60 57.11
Adapter 13.29 (4.66%) 78.46 76.38 77.36 67.12 80.33 82.51 67.77 71.71 69.58 71.24 74.25
READ 0.14 (0.05%) 88.30 85.15 89.76 75.80 86.69 86.62 74.99 82.38 84.65 79.60 83.39
Table 2: TLG results on the TVSum dataset. We report the mean average precision (mAP) and the number of trainable parameters (#params). means the gain of READ is statistically significant at the 0.05 level.
Method #params (M) mAP
Full 15.88 (100%) 36.14
Bias 0.06 (0.38%) 24.89
Partial 1.05 (6.61%) 26.37
Proj 7.31 (46.03%) 32.71
LoRA 0.19 (1.20%) 33.96
Prompt 0.04 (0.25%) 25.86
Adapter 0.20 (1.26%) 33.61
READ 0.19 (1.20%) 36.74
Table 3: TLG results on the QVHighlights dataset. We report the mean average precision (mAP) and the number of trainable parameters (#params). means the gain of READ is statistically significant at the 0.05 level.
We conduct extensive experiments to evaluate the effectiveness of our READ framework. We first describe the experimental settings, covering the downstream tasks, evaluation metrics, pre-trained backbones, baseline approaches, and implementation details. We then present the numerical results of our method with baseline models, then provide ablation study and thorough analysis to explore various configurations. Eventually, we perform qualitative assessments to further elucidate the behavior of our framework.
Experimental Settings
@@
Method #params (M) R1 R2 RL
Full 249.67 (100%) 35.72 11.88 30.00
Bias 0.20 (0.08%) 30.51 8.20 23.00
Partial 16.54 (6.62%) 31.55 8.63 13.65
Proj 38.60 (15.46%) 32.76 9.11 30.01
LoRA 1.19 (0.48%) 40.04 20.36 35.98
Prompt 0.05 (0.02%) 31.55 8.63 14.90
Adapter 1.20 (0.48%) 41.52 20.75 36.88
READ 1.17 (0.47%) 44.01 21.91 37.91
Table 4: VLS results on the How2 dataset with the VG-BART model. We report the ROUGE-1, ROUGE-2, and ROUGE-L scores, with the number of trainable parameters (#params). means the gain of READ is statistically significant at the 0.05 level.
Method #params (M) R1 R2 RL
Full 333.16 (100%) 32.37 8.07 26.53
Bias 0.07 (0.02%) 27.03 4.53 19.51
Partial 16.52 (4.96%) 27.83 4.92 10.49
Proj 24.67 (7.40%) 29.16 5.50 26.76
LoRA 1.93 (0.58%) 36.63 16.72 32.77
Prompt 0.18 (0.05%) 28.33 5.12 11.75
Adapter 1.95 (0.59%) 37.83 17.45 33.87
READ 1.91 (0.57%) 40.12 18.71 34.42
Table 5: VLS results on the How2 dataset with the VG-T5 model. We report the ROUGE-1, ROUGE-2, and ROUGE-L scores, with the number of trainable parameters (#params). means the gain of READ is statistically significant at the 0.05 level.
Method mAP - YouTube Highlights R2 - How2
No VLA 73.80 18.22
VLA 74.41 20.01
PVLA 76.12 21.91
Table 6: Partial video-language alignment (PVLA) ablation experiments on YouTube Highlights and How2. We color the settings we implement for our READ method.
Video Language query POT distance AP
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Girl’s expression after getting proposal 27.17 90.28
A boy showing his arm after being stung at the beach 56.38 30.00
Table 7: Case study on the temporal language grounding benchmark. We extract the POT distance between video and language of two inputs with different language queries and measure the respective AP performance change.
Recurrent architecture #params - UMT (M) mAP - YouTube Highlights #params - VG-BART (M) R2 - How2
GRU 0.16 76.07 1.17 21.85
LSTM 0.16 76.08 1.17 21.87
RNN 0.16 76.12 1.17 21.91
Table 8: Recurrent architecture ablation experiment on YouTube Highlights and How2. We color the settings we implement for our READ method.
Distance method mAP - YouTube Highlights R2 - How2
AvgPool - Cosine 71.65 20.32
MaxPool - Cosine 74.37 21.36
AvgPool - L2 72.11 20.73
MaxPool - L2 74.39 21.02
Partial OT 76.12 21.91
Table 9: Distance method ablation experiments on YouTube Highlights and How2. We color the settings we implement for our READ method.
@@
Main Results
For main comparison of our READ with baseline methods, we denote the results of YouTube Highlights in Table 1, TVSum in Table 2, QVHighlights in Table 3, and How2 in Tables 4 and 5.
Temporal language grounding (TLG). For the YouTube Highlights dataset, our READ framework substantially outperforms the Full fine-tuning approach (e.g. 1.36% on average, 1.75% in the Dog domain, and 1.21% in the Surfing domain), while updating far less parameters (0.16M vs. 283.97M). We significantly surpass all other efficient fine-tuning methods as well, e.g. with an improvement of 10.96% over the Adapter in the Gym category, or 5.78% over LoRA in the Skating category.
@@
These results demonstrate that our READ framework can efficiently model video-language inputs to polish the low-resource temporal language grounding performance of various pre-trained Transformer models.
@@
Ablation Studies
We ablate our READ framework to discover what factors result in the demonstrated efficiency and observe several intriguing properties. Our ablation studies are all conducted on the YouTube Highlights and How2 test set.